OpenStudy (anonymous):

Show all the work to solve |x+3| +7 ≤ 10

OpenStudy (anonymous):

CaseI X+3 + 7 < = 10 X < = 0 CaseII -x - 3 + 7 < =10 -x + 4 <=10 -x < = 6 x> = -6

OpenStudy (anonymous):

nw combine.... -6 <= x <= 0

OpenStudy (anonymous):

THANK YOU!!